OpenStudy (anonymous):

What is the missing number? –3.4 + (–8 + 2) = [ + (–8)] + 2

OpenStudy (lalipapa123):

where the ................... i.e, blank?

OpenStudy (anonymous):

why are so many poeple here

OpenStudy (puppylover52):

because we want to help

OpenStudy (anonymous):

ITS THE ONLY OTHER NUMBER WTH A NEGATIVE oops caps sorry

OpenStudy (lalipapa123):

do u have any idea about this?

OpenStudy (igreen):

\(\sf –3.4 + (–8 + 2) = [\_\_ + (–8)] + 2\) THis is the question?

OpenStudy (anonymous):

no that's y I asked you guys

OpenStudy (igreen):

Okay, first simplify the left side. -8 + 2 = ?

OpenStudy (igreen):

@lalipapa123 Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. Please read the \(\bf\color{blue}{\underline{{\href{http:///openstudy.com/code-of-conduct}{Code-of-Conduct}}}}\) to learn more.

OpenStudy (anonymous):

-6

OpenStudy (lalipapa123):

kkk

OpenStudy (lalipapa123):

now solve the righht side numbers

OpenStudy (anonymous):

-6 iGreen

OpenStudy (anonymous):

-8 + 2 = -6

OpenStudy (anonymous):

is it -3.4 the answer

OpenStudy (lalipapa123):

–3.4 + (–8 + 2) = [ + (–8)] + 2 -3.4+(-6) = [ +(-8)]+2 then you will get -9.4 = -6 now let the left side come to right side -9.4 + 6 =3.4

OpenStudy (anonymous):

you mean -3.4

OpenStudy (lalipapa123):

yes!

OpenStudy (anonymous):

thanks you guys give igreen one I give you one

OpenStudy (lalipapa123):

gave!

OpenStudy (puppylover52):

ok no problem

OpenStudy (anonymous):

who gave iGreen one

OpenStudy (lalipapa123):

i

OpenStudy (anonymous):

there

OpenStudy (lalipapa123):

thanks!

OpenStudy (lalipapa123):

anyother questions?

OpenStudy (anonymous):

can you guys help me on my quistion

OpenStudy (lalipapa123):

yeah sure

OpenStudy (puppylover52):

sure

OpenStudy (anonymous):

I called you guys

OpenStudy (lalipapa123):

where?

OpenStudy (igreen):

Yes..now we add -3.4 to it.. @coolman500

OpenStudy (anonymous):

ok that was the answer thanks